Please can you help me with this equation?
x^2/3 - 8x^1/3+15=0

Answer by stanbon(48575)

x^2/3 - 8x^1/3+15=0
-------------
Let w = x^(1/3)
Substitute to get:
w^2-8w+15 = 0
Factor:
(w-5)(x-3)=0
w=5 or w=3
--------------
solve for "x":
x^(1/3)= 5
x = 5^3= 125
----------
x^(1/3)=3
x = 3^3 = 27
------------
Final Answer: x = 125 or x=27
=================
Cheers,
Stan H.